Calculate the value of a bond that matures in 12 years and has a $1,000 par value. The annual coupon interest rate is 14

percent and the​ market's required yield to maturity on a​ comparable-risk bond is 12 percent.

The value of the bond is_____?
Answer #1

Annual coupon=1000*14%=140

Hence value of bond=Annual coupon*Present value of annuity factor(12%,12)+$1000*Present value of discounting factor(12%,12)

=140*6.194374225+$1000*0.256675092

=$1123.89(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=140[1-(1.12)^-12]/0.12

=140*6.194374225

2.Present value of discounting factor=1000/1.12^12

=1000*0.256675092